Solve the triangle, if possible. Determine the number of possible solutions. A = 82.3° a = 11.5 b = 9.7
Answer #1

given A-82.3°, a 115, 6-9.7 a ² = 6²762_2 bccosa a (11.57-19-7)+04-2195301005823) 132.25-94.09+02_19.4(0) (0.1339) c?_2.50 +9с. - bTVь час 24 с. -f N 2.0* 4x-38-k) 2С, )c=2.55 C & -5.05 .:C=7,55 Crow a²+C²_6² COSBE 2 ac 111.77.5) 9.73 2(11.5) (25) 132.257 56.25-94.09 172.5 94.40 172.5 - 5 6!9441 COSB3 172,5 COSB = 0.54 .:B - 56.839 Cosc = a + b²c2 zab =(11.5)²+lan=0.02 2 (11.5) (9.7) = 132.25 +94.09-56.25 223.1 22